OpenStudy (anonymous):

Solve. Even though you might find the answer quickly in some other way, practice using the five-step problem-solving process in order to build the skill of problem solving. Budget overruns. The massive roadworks project in Boston known as The Big Dig cost approximately $14.6 billion. This cost was 484% more than the original estimate. What was the original estimate of the cost of The Big Dig?

OpenStudy (mertsj):

Let x = original cost 4.84x = 14.6 billion 4.84x/4.84 = 14.6/4.84 x = 3.02 billion

OpenStudy (whpalmer4):

That's close, but incorrect :-) The key is that the cost was 484% more than the original estimate, which is not the same as 484% of the original estimate. Original estimate*(1+percentage more/100) = Final You can see this to be true if you consider some obvious percentages: if the cost is 0% more than the original estimate, the formula reduces to original*(1+0/100) = original; if the cost is 100% more than the original estimate, you get original*(1+100/100) = 2 * original. so, in this case, x = original cost. final cost=(1+484/100)*original or $14.6 billion = (1+4.84)*x and so x = $14.6 billion / 5.84 = $2.5 billion The numbers I've seen say the original estimate was $2.6 billion, not 2.5, but what's $10^8 between friends? :-)

OpenStudy (anonymous):

Let x be the original estimate. Solve the following for x:\[100\%x+484\% x \text{=} 14.6*10^9\]\[1.00x+4.84 x = 14.6*10^9\]\[x+4.84 x = 14.6*10^9 \]\[x=2.5\times 10^9 \]The above was solved with the aid of Mathematica. I don't want to ruffle any feathers here, but in some circles, "between friends" would be read, between democrats when it comes to that project.